a body is weighing 20 kg has a velocity of 5m/s, after one second of starting from the mean position. if the time period is 6 seconds,find the k.E.and P.E. and total energy.
1
Expert's answer
2013-03-14T14:15:20-0400
a body is weighing 20 kg has a velocity of 5m/s, after one second of starting from the mean position. if the time period is 6 seconds, find the k.E.and P.E. and total energy.


The acceleration of a body is


a = 5[m/s]/1[s] = 5[m/s²].


6 seconds after starting velocity of a body becomes


V = a·6[s] = 5[m/s²]·6[s] = 30[m/s].


So, kinetic energy of a body is


K = mV²/2 = 20[kg]·(30[m/s])²/2 = 9000[J].


We have nothing to say about potential energy 'cause we have no data on the force field a body to be in.